An option is sold short (call and put). The required margin is $20,000. The cost of...

  1. An option is sold short (call and put). The required margin is $20,000. The cost of margin debt is 8%/ year. The position is held for 45 days. What is the cost of margin?
Answer #1

Cost of Margin = Margin* Cost of Margin*45/365 = 20,000 * 8%*45/365 = 197.26

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well